What would y= -1/2x + 3 be written in standard form?
Found 2 solutions by TimothyLamb, jim_thompson5910:
Answer by TimothyLamb(4379)

y = -1/2x + 3
Answer:
(1/2)x + (1)y = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by jim_thompson5910(35256)

y= -1/2x + 3

2y= -x + 6 ... multiply every term by 2 to clear out the fraction

2y + x = 6

x + 2y = 6

=======================================================
Final Answer:

x + 2y = 6